For an event if you purchased 8 adult tickets and 6 children tickets for your party at a cost of 132 dollars.A friend purchased 10 adult tickets and 4 children tickets for 144 dollars.Find the price of the adult ticket and the price of the children ticket.
Answer by mananth(16946)

x= cost of adult ticket
y= cost of child ticket

8.00 x + 6.00 y = 132.00 .............1
Total value
10.00 x + 4.00 y = 144.00 .............2
Eliminate y
multiply (1)by -4.00
Multiply (2) by 6.00
-32.00 x -24.00 y = -528.00
60.00 x + 24.00 y = 864.00
Add the two equations
28.00 x = 336.00
/ 28.00
x = 12.00
plug value of x in (1)
8.00 x + 6.00 y = 132.00
96.00 + 6.00 y = 132.00
6.00 y = 132.00 -96.00
6.00 y = 36.00
y = 6.00
x= 12 cost of adult ticket $
y= 6 cost of child ticket $
m.ananth@hotmail.ca